Help Periodic Inventory Methods/Valuation Hyper Company had a beginning inventory on January 1 of 160 units...

Help

Periodic Inventory Methods/Valuation Hyper Company had a beginning inventory on January 1 of 160 units of Product 4-18-19 at a cost of $20 per unit. During the year, the following purchases were made.

Mar. 15 400 units at $23

Sept. 4 330 units at $26

July 20 250 units at $24

Dec. 2 100 units at $29

1,000 units were sold.

Hyper Company uses a periodic inventory system. Instructions

(a) Determine the cost of goods available for sale.

(b) Determine (1) the ending inventory, and (2) the cost of goods sold under each of the assumed cost flow methods (FIFO, LIFO, and average-cost). (c) Which cost flow method results in (1) the highest inventory amount for the balance sheet, and (2) the highest cost of goods sold for the income statement?
Answer #1

a) Calculate cost of goods available for sale

Unit Unit cost Total Cost
Beginning inventory 160 20 3200
Purchase Mar 15 400 23 9200
Purchase Sept 4 330 26 8580
Purchase July 20 250 24 6000
Purchase Dec 2 100 29 2900
Total 1240 29880

Cost of goods available for Sales = $29880

2) Calculate following

FIFO LIFO Average cost
Ending inventory (100*29+140*24) = 6260 (160*20+80*23) = 5040 29880/1240*240 = 5783
Cost of goods sold 29880-6260 = 23620 29880-5040 = 24840 29880-5783 = 24097

c) Highest Inventory amount = FIFO = $6260

d) Highes cost of goods sold = LIFO = $24840